OpenStudy (abbycross167):

can someone please help me with this math question? please? I'll fan and give a medal... please show me how to graph -1/2 + (1/2)y >= x

OpenStudy (abbycross167):

What do you mean? @agent0smith that is the whole question... I just need to show the steps to graph that equation

OpenStudy (agent0smith):

First get it into y = mx+b form

OpenStudy (abbycross167):

y = -1/2x+(1/2) like this?

OpenStudy (agent0smith):

No...

OpenStudy (agent0smith):

Add 1/2 to both sides you did correct. But then you have (1/2)y = x+(1/2)

OpenStudy (abbycross167):

oh okay so, the y = mx+b form of this- -1/2 + (1/2)y >= x = (1/2)y = x+(1/2)

OpenStudy (agent0smith):

Then how do you get rid of the 1/2 that is being multiplied by y?

OpenStudy (abbycross167):

umm i'm not sure

OpenStudy (agent0smith):

Multiply everything by 2/1

OpenStudy (abbycross167):

ok so - 2*1(1/2)y = x+(1/2)*2*1

OpenStudy (agent0smith):

multiply EVERY term by 2/1

OpenStudy (agent0smith):

2/1 not 2*1...

OpenStudy (abbycross167):

2/1(1/2)y = x*2/1+(1/2)*2/1

OpenStudy (agent0smith):

Good.

OpenStudy (abbycross167):

then what do I do?

OpenStudy (agent0smith):

Do math. Simplify.

OpenStudy (abbycross167):

2/1*(1/2)y = x*2/1+(1/2)*2/1 2/1*(1/2) = 1 2/1+(1/2)*2/1 = 3

OpenStudy (abbycross167):

y = 3x+1? or would it be 2x?

OpenStudy (agent0smith):

No idea what you did :/

OpenStudy (agent0smith):

\[\large \frac{ 2 }{ 1 }*\frac{ 1 }{ 2 }y=x*\frac{ 2 }{1 }+\frac{ 1 }{2 }*\frac{ 2 }{ 1 }\]

OpenStudy (agent0smith):

Yes, 2/1*1/2 is 1. How did you get a 3??

OpenStudy (abbycross167):

I multiplied...

OpenStudy (agent0smith):

Keep in mind all three terms are separate. You CANNOT add any of them together.

OpenStudy (abbycross167):

oh okay

OpenStudy (abbycross167):

so 2/1+1/2*2/1= 2 ?

OpenStudy (agent0smith):

Why are you adding anything together at all? You should only be simplifying with multiplying... the x and 2/1 are attached by multiplication

OpenStudy (abbycross167):

2x+1?

OpenStudy (agent0smith):

Much better

OpenStudy (agent0smith):

now graph y=2x+1

OpenStudy (abbycross167):

thank you very much :) I really appreciate your help!!

OpenStudy (agent0smith):

Welcome :)

OpenStudy (agent0smith):

Don't forget you're graphing the inequality \[\large y \ge 2x +1\]